OpenStudy (klepez):

Which of the quadratic functions has the narrowest graph? (1 point) y = –x^2 y = x^2 y = 4x^2 y = x^2

OpenStudy (anonymous):

two function are same (y = x^2)

OpenStudy (anonymous):

@klepez two function are repeated

OpenStudy (klepez):

Which of the quadratic functions has the narrowest graph? (1 point) y = –x2 y = 1/4x^2 y = 4x2 y = 1/9x^2 thats what its soposed to look like sorry.

OpenStudy (klepez):

thank you!

OpenStudy (anonymous):

@klepez welcome^-|-^